OpenStudy (anonymous):

Rationalize the denominator square root of 20/x

OpenStudy (agreene):

are you asking \(\large\sqrt{\frac{20}{x}}\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i have the answer, i just don't know how to get it

OpenStudy (agreene):

Well, I was going to say I dont know what exactly they're wanting, because it doesnt really make sense...

OpenStudy (agreene):

\[2\sqrt{5}\sqrt{x^{-1}}\] is how I would answer, but its not "Rationalizing the denominator"

OpenStudy (anonymous):

\[\frac{ 2\sqrt{5x} }{ x }\] is the answer

OpenStudy (anonymous):

i just don't get why x is in the numerator

OpenStudy (agreene):

I got there like this: \[\sqrt{\frac{20}{x}} = \sqrt{20}\sqrt{-1} = 2\sqrt{5} \sqrt{x^{-1}}\]

OpenStudy (agreene):

Well, I suppose it's true if x is always positive... but it's technically wrong.

OpenStudy (anonymous):

thank you for your help!

OpenStudy (agreene):

\[\sqrt{\frac{20}{x}}=\frac{\sqrt{20}}{\sqrt{x}} = \frac{2\sqrt{5}}{\sqrt{x}}\]

OpenStudy (anonymous):

YESSS that has to be right.

OpenStudy (agreene):

that x with the 5 is to keep it from being 0/0

OpenStudy (agreene):

but technically, they arent the same.

OpenStudy (anonymous):

so does the x have to be written?

OpenStudy (agreene):

I mean, if you're instructor requires it... lol

OpenStudy (anonymous):

If x is not a perfect square the denominator will be irrational, so multiplying and dividing by a factor of square root of x will rationalize the denominator

OpenStudy (anonymous):

Feels rather unnecessary though

OpenStudy (triciaal):